# COSMOLOGICAL DYNAMICS AND OBSERVATIONAL CONSTRAINTS ON A VIABLE $f(Q)$ NON-METRIC GRAVITY MODEL

### 19 Jan 2024 v2

@@ Abstract
keywords:
Modified gravity; Dark energy; $f(Q)$ gravity; Parameter constraints.
[…]
PACS numbers: 04.50.Kd, 98.80.-k
@@ 1 Introduction
For over two decades, significant efforts have been dedicated to cosmological research in pursuit of an explanation of the observed late-time cosmic acceleration. The primary avenue of investigation involves the introduction of a novel energy component within the Universe, termed dark energy (DE), which is distinguished by its negative pressure. Nevertheless, as of the present time, a definitive and satisfactory resolution to the enigma of DE remains elusive; its incorporation into the framework of fundamental physics theories continues to challenge researchers (for a comprehensive review about this topic see Refs. 1, 2, 3).
One of the intriguing approaches for elucidating the late-time cosmic acceleration, beyond the inclusion of either DE or novel forms of matter to account for this phenomenon, lies within the field of modified gravity theories (see e.g. Refs. 4, 5, 6, 7, 8, 9 for a review). Usually, in this framework, the fundamental action is built assuming generalized functions of the scalar curvature (the so-called $f(R)$ theories), general higher-order theories, scalar–tensor theories of gravitation, etc. Recently, a new proposal has emerged within the realm of modified theories of gravitation. These particular theories, in which gravitational interactions are governed by non-metricity, with curvature and torsion being rendered negligible, are known as $f(Q)$ theories or $f(Q)$ symmetric teleparallel gravity, where $Q$ is the non-metricity scalar [10, 11, 12, 13, 14, 15]. These theoretical frameworks hold the potential to provide fresh perspectives on the phenomenon of cosmic acceleration, stemming from the inherent consequences of an alternative geometry as opposed to the conventional Riemannian framework (see Ref. 16 for a recent and extensive review on this topic).
@@
For instance, in Ref. 17 the authors performed an observational analysis of several modified $f(Q)$ models using the redshift approach, where the $f(Q)$ Lagrangian is reformulated as an explicit function of the redshift, $f(z)$. Various different polynomial parameterizations of $f(z)$ are proposed, including new terms which would allow for deviations from the $\Lambda$CDM model. In Ref. 31 a new parametrization of the Hubble parameter is proposed in a model-independent way and apply it to the Friedmann equations in the FLRW Universe. Also, the authors of Ref. 43 implemented a parametrization scheme for the Hubble parameter, obtaining an exact solution for the field equations in $f(Q)$ cosmology.
The proposal presented here is a new viable $f(Q)$ gravity model which can be represented as a perturbation of the $\Lambda$CDM model. This model is inspired by an exponential $f(R)$ gravity model studied recently in the literature [52, 53, 54], which depends upon two parameters, $b$ and $n$. Furthermore, instead of using the aforementioned strategies, we have chosen to implement the formalism developed in Ref. 55 within the context of $f(R)$ gravity. This formalism involves expressing the solution to the field equations, i.e., $H^{2}(N)$, as a perturbative series expansion in terms of the deviation parameter $b$ (making $n=1$), with the $\Lambda$CDM model solution as a reference. Additionally, we deduce approximated analytical expressions for the parameters $w_{\rm{DE}}$, $w_{\rm{eff}}$, and $\Omega_{\rm{DE}}$, as well as the deceleration parameter $q$ in terms of the redshift $z$. Finally, in order to verify the viability of this approximate analytical solution for $H(z)$ in late-times, we examine the behavior of these cosmological parameters, and perform a fit to observational data of the Hubble parameter ($H(z)_{{\rm obs}}$) and the apparent magnitude ($m_{b}(z)_{{\rm obs}}$, Pantheon SN sample) in terms of the redshift, to obtain constraints on the model parameter $b$, together with the present values of the expansion rate, $H_{0}$, and the matter density parameter, $\Omega_{m0}$.
@@ 2 The $f(Q)$ gravity: a brief review
$S=\int{d^{4}x\sqrt{-g}\left(\frac{f(Q)}{2\kappa^{2}}+\mathcal{L}_{\rm{M}}\right)},$ (1)
@@
$Q_{\gamma\mu\nu}\equiv\nabla_{\gamma}g_{\mu\nu}=\partial_{\gamma}g_{\mu\nu}-\Gamma^{\beta}_{\mkern 10.0mu\gamma\mu}g_{\beta\nu}-\Gamma^{\beta}_{\mkern 10.0mu\gamma\nu}g_{\beta\mu},$ (2)
@@
$\Gamma^{\lambda}\\ _{\mu\nu}=\left\{{}^{\lambda}\\ _{\mu\nu}\right\}+K^{\lambda}\\ {}_{\mu\nu}+L^{\lambda}\\ _{\mu\nu},$ (3)
@@
{}$\left\{{}^{\lambda}\\ _{\mu\nu}\right\}\equiv\frac{1}{2}g^{\lambda\beta}\left(\partial_{\mu}g_{\beta\nu}+\partial_{\nu}g_{\beta\mu}-\partial_{\beta}g_{\mu\nu}\right);$ (4)
@@
$K^{\lambda}\\ _{\mu\nu}\equiv\frac{1}{2}T^{\lambda}\\ _{\mu\nu}+T^{\,\,\,\,\,\lambda}_{(\mu\,\,\,\,\nu)},$ (5)
@@
$L^{\beta}_{\mkern 10.0mu\mu\nu}=\frac{1}{2}Q^{\beta}_{\mkern 10.0mu\mu\nu}-Q_{(\mu\nu)}^{\mkern 35.0mu\beta}.$ (6)
@@
$P^{\beta}_{\mkern 10.0mu\mu\nu}=-\frac{1}{2}L^{\beta}_{\mkern 10.0mu\mu\nu}+\frac{1}{4}(Q^{\beta}-\tilde{Q}^{\beta})g_{\mu\nu}-\frac{1}{4}\delta^{\beta}_{\,\,\,(\mu}Q_{\nu)},$ (8)
the other quantities present in Eq. (8) are given by
$Q_{\beta}=g^{\mu\nu}Q_{\beta\mu\nu},\quad\tilde{Q}_{\beta}=g^{\mu\nu}Q_{\mu\beta\nu}.$ (9)
Variation of the action (1) with respect to the metric gives the equation of motion [12, 13],
$\frac{2}{\sqrt{-g}}\nabla_{\beta}(f_{Q}\sqrt{-g}P^{\beta}_{\mkern 10.0mu\mu\nu})+\frac{1}{2}fg_{\mu\nu}+f_{Q}(P_{\mu\beta\lambda}Q_{\nu}^{\mkern 10.0mu\beta\lambda}-2Q_{\beta\lambda\mu}P^{\beta\lambda}\mkern 1.0mu\nu)=-T_{\mu\nu},$ (10)
@@
$T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta\sqrt{-g}\mathcal{L}_{\rm{M}}}{\delta g^{\mu\nu}},$ (11)
@@
$ds^{2}=-dt^{2}+a(t)^{2}\delta_{ij}dx^{i}dx^{j},$ (12)
with $a(t)$ representing the scale factor, the time and spatial components of Eq. (10) are given, respectively, by [12, 13]
$6f_{Q}H^{2}-\frac{1}{2}f=\rho_{\rm{m}}+\rho_{\rm{r}},$ (13)
@@
$(12H^{2}f_{QQ}+f_{Q})\dot{H}=-\frac{1}{2}\left(\rho_{\rm{m}}+\frac{4}{3}\rho_{\rm{r}}\right),$ (14)
where $\rho_{\rm{m}}$ is the matter density and $\rho_{\rm{r}}$ denotes the density of radiation. $f_{QQ}=\frac{d^{2}f}{dQ^{2}}$, the over-dot denotes a derivative with respect to the cosmic time $t$ and $H\equiv\dot{a}/a$ is the Hubble parameter. Additionally, we have considered $\kappa^{2}=1$. Further, the non-metricity scalar associated with the metric (12) is given by
$Q=6H^{2}.$ (15)
@@
$\dot{\rho}_{\rm{m}}+3H\rho_{\rm{m}}=0,\quad\dot{\rho}_{\rm{r}}+4H\rho_{\rm{r}}=0.$ (16)
As usual in the literature, it is possible to rewrite the field equations (13) and (14) in the Einstein-Hilbert form:
$3H^{2}=\rho,$ (17)
$-2\dot{H}^{2}=\rho+p,$ (18)
where $\rho=\rho_{\rm{m}}+\rho_{\rm{r}}+\rho_{\rm{DE}}$ and $p=p_{\rm{m}}+p_{\rm{r}}+p_{\rm{DE}}$ correspond to the total effective energy pressure densities of the cosmological fluid, respectively. In this case, the dark energy component has a geometric origin, and after some manipulation in Eqs. (13) and (14), we obtain the effective dark energy and pressure corresponding to the $f(Q)$-theory given by
$\rho_{\rm{DE}}=\frac{1}{2}f+3H^{2}(1-2f_{Q}),$ (19)
@@
$w_{\rm{DE}}=\frac{p_{\rm{DE}}}{\rho_{\rm{DE}}}=-1+\frac{2\dot{H}(12H^{2}f_{QQ}+f_{Q}-1)}{\frac{1}{2}f+3H^{2}(1-2f_{Q})}.$ (22)
@@ 3 Cosmological dynamics in late-time
$f(Q)=Q+2\Lambda e^{-\left(\frac{b\Lambda}{Q}\right)^{n}},$ (23)
where $\Lambda$ is the cosmological constant, and $b$ and $n$ are real dimensionless parameters. This model is inspired in that studied in Refs. 52, 53, 54, but in the context of $f(R)$ gravity. It is evident that for $b=0$ the model given by Eq. (23) is equivalent to GR plus the cosmological constant. In particular, from the structure of this model, it can be seen as a perturbative deviation from the $\Lambda$CDM Lagrangian. In this sense, this model can be arbitrarily close to $\Lambda$CDM, depending on the parameter $b$, i.e., the exponential form presented in the Eq. (23) elegantly and succinctly encapsulates the perturbative nature of the model across all orders in the parameter $b$, which is evident from a Taylor series expansion about this parameter,
$f(Q)=(Q+2\Lambda)-\frac{2\Lambda^{2}b}{Q}+\frac{\Lambda^{3}b^{2}}{Q^{2}}-\frac{\Lambda^{4}b^{3}}{3Q^{3}}+\frac{\Lambda^{5}b^{4}}{12Q^{4}}-\cdots,$ (24)
where for simplicity we have considered $n=1$. It should be highlighted that in the literature also other exponential $f(Q)$ gravity models have been intensively studied (see e.g., Refs. 22, 28, 29, 30, 36, 37, 38, 41).
Following the procedure carried out in Ref. 55, we rewrite Eq. (13) in terms of $N=\ln{a}$
$2f_{Q}H^{2}(N)-\frac{1}{6}f-(\Omega_{m0}e^{-3N}+\Omega_{r0}e^{-4N})H_{0}^{2}=0,$ (25)
where we have introduced the current energy density parameters, $\Omega_{m0}=\frac{\rho_{m0}}{3M_{\rm{p}}^{2}H_{0}^{2}}$ and $\Omega_{r0}=\frac{\rho_{r0}}{3M_{\rm{p}}^{2}H_{0}^{2}}$; here $M_{\rm{p}}=1$, and $H_{0}$ corresponds to the current Hubble parameter. Implicitly, we consider by simplicity $n=1$ and $\Lambda=3(1-\Omega_{m0}-\Omega_{r0})H_{0}^{2}$. Replacing Eq. (23) in Eq. (25) and taking into account Eq. (15), we obtain an algebraic equation for $H^{2}(N)$, given by
$H^{2}(N)-(\Omega_{m0}e^{-3N}+\Omega_{r0}e^{-4N})H_{0}^{2}$ (26)
@@
Since the above is a transcendental equation, it cannot be solved for $H^{2}(N)$ in explicit form. Another possibility involves constructing a first-order differential equation for $H^{2}(N)$ using Eqs. (13) and (14). However, in this scenario the equation is nonlinear and lacks an analytical solution for $H^{2}(N)$.
Usually, in the literature for circumventing this issue, a common practice is to propose a parametrization of the Hubble parameter in terms of the redshift [17, 31, 43], among other strategies. In this work instead of use that strategy, we use the formalism developed in Ref. 55 in the context of $f(R)$ gravity, which consists in expressing the solution to Eq. (25), i.e., $H^{2}(N)$, as a perturbative series expansion in terms of the deviation parameter $b$, with respect to the $\Lambda$CDM model solution as a reference, as follows
$H^{2}(N)=H^{2}_{\Lambda}(N)+\sum_{i=1}^{M}b^{i}\delta H_{i}^{2}(N),$ (27)
@@
$\frac{H^{2}_{\Lambda}(N)}{H_{0}^{2}}=\Omega_{m0}\,e^{-3N}+\Omega_{r0}\,e^{-4N}+(1-\Omega_{m0}-\Omega_{r0})=E_{\Lambda}^{2}(N),$ (28)
further, as has been demonstrate in Ref. 55, we can consider only two terms in the above series expansion; in doing so, Eq. (27) reduces to
$H^{2}(N)\approx H^{2}_{\Lambda}(N)+b\,\delta H_{1}^{2}(N)+b^{2}\delta H_{2}^{2}(N).$ (29)
Replacing Eq. (29) in Eq. (26) and considering up to second order in $b$, it is straightforward to calculate $\delta H_{1}^{2}(N)$ and $\delta H_{2}^{2}(N)$, which are given by
$\frac{\delta H_{1}^{2}(N)}{H_{0}^{2}}=-\frac{3H_{0}^{2}(\Omega_{m0}+\Omega_{r0}-1)^{2}}{2H_{\Lambda}^{2}(N)},$ (30)
$\frac{\delta H_{2}^{2}(N)}{H_{0}^{2}}=-\frac{H_{0}^{4}(\Omega_{m0}+\Omega_{r0}-1)^{3}[18H_{0}^{2}(\Omega_{m0}+\Omega_{r0}-1)+5H_{\Lambda}^{2}(N)]}{8(H_{\Lambda}^{2}(N))^{3}}.$ (31)
Now, replacing Eqs. (30) and (31) in Eq. (29), and using Eq. (28), we obtain an approximate solution for the Hubble parameter $H(z)$:
$E^{2}(z)=$ $\frac{H^{2}(z)}{H_{0}^{2}}\approx 1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}-\frac{3(\Omega_{m0}-1)^{2}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]}b$ (32)
@@
$w_{\rm{DE}}=-1+\frac{1}{3}(1+z)\frac{(\rho_{\rm{DE}}(z))^{\prime}}{\rho_{\rm{DE}}(z)},$ (33)
$w_{\rm{eff}}=-1+\frac{1}{3}(1+z)\frac{(E^{2}(z))^{\prime}}{E^{2}(z)},$ (34)
$\Omega_{\rm{DE}}=1-\frac{\Omega_{m0}(1+z)^{3}}{E^{2}(z)};$ (35)
@@
$q=-1+\frac{1}{2}(1+z)\frac{(E^{2}(z))^{\prime}}{E^{2}(z)},$ (36)
where the prime denotes differentiation with respect to $z$. Using Eqs. (19) and (32) and considering up to second order expansion in $b$, we obtain approximated analytical expressions for the above parameters in terms of the redshift $z$, as follows:
$w_{\rm{DE}}\approx$ $-1-\frac{3[(1+z)^{3}(\Omega_{m0}-1)\Omega_{m0}]}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{2}}b$ (37)
@@
$w_{\rm{eff}}\approx$ $-1+\frac{(1+z)^{3}\Omega_{m0}}{1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}}+\frac{3(1+z)^{3}(\Omega_{m0}-1)^{2}\Omega_{m0}}{[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{3}}b$ (38)
@@
$\Omega_{\rm{DE}}\approx$ $1-\frac{(1+z)^{3}\Omega_{m0}}{1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}}-\frac{3(1+z)^{3}(1-\Omega_{m0})^{2}\Omega_{m0}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{3}}b$ (39)
@@
$q\approx$ $-1+\frac{3(1+z)^{3}\Omega_{m0}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]}+\frac{9(1+z)^{3}(\Omega_{m0}-1)^{2}\Omega_{m0}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{3}}b$ (40)
@@
With Eqs. (37)-(40), we can plot the evolution of each parameter in terms of the redshift $z$. Additionally, in order to compare the results with the $\Lambda$CDM model, we have also incorporated in the corresponding plots the behavior associated with each quantity defined by Eqs. (33)-(36), but using Eq. (28) instead of (32).

### 30 Jan 2024 v3

@@ Abstract
Keywords: Modified gravity; Dark energy; $f(Q)$ gravity; Parameter constraints.
PACS: 04.50.Kd, 98.80.-k
@@ 1 Introduction
For over two decades, significant efforts have been dedicated to cosmological research in pursuit of an explanation of the observed late-time cosmic acceleration. The primary avenue of investigation involves the introduction of a novel energy component within the Universe, termed dark energy (DE), which is distinguished by its negative pressure. Nevertheless, as of the present time, a definitive and satisfactory resolution to the enigma of DE remains elusive; its incorporation into the framework of fundamental physics theories continues to challenge researchers (for a comprehensive review about this topic see Refs. [1, 2, 3]).
One of the intriguing approaches for elucidating the late-time cosmic acceleration, beyond the inclusion of either DE or novel forms of matter to account for this phenomenon, lies within the field of modified gravity theories (see e.g. Refs. [4, 5, 6, 7, 8, 9] for a review). Usually, in this framework, the fundamental action is built assuming generalized functions of the scalar curvature (the so-called $f(R)$ theories), general higher-order theories, scalar–tensor theories of gravitation, etc. Recently, a new proposal has emerged within the realm of modified theories of gravitation. These particular theories, in which gravitational interactions are governed by non-metricity, with curvature and torsion being rendered negligible, are known as $f(Q)$ theories or $f(Q)$ symmetric teleparallel gravity, where $Q$ is the non-metricity scalar [10, 11, 12, 13, 14, 15]. These theoretical frameworks hold the potential to provide fresh perspectives on the phenomenon of cosmic acceleration, stemming from the inherent consequences of an alternative geometry as opposed to the conventional Riemannian framework (see Ref. [16] for a recent and extensive review on this topic).
@@
For instance, in Ref. [17] the authors performed an observational analysis of several modified $f(Q)$ models using the redshift approach, where the $f(Q)$ Lagrangian is reformulated as an explicit function of the redshift, $f(z)$. Various different polynomial parameterizations of $f(z)$ are proposed, including new terms which would allow for deviations from the $\Lambda$CDM model. In Ref. [31] a new parametrization of the Hubble parameter is proposed in a model-independent way and apply it to the Friedmann equations in the FLRW Universe. Also, the authors of Ref. [43] implemented a parametrization scheme for the Hubble parameter, obtaining an exact solution for the field equations in $f(Q)$ cosmology.
The proposal presented here is a new viable $f(Q)$ gravity model which can be represented as a perturbation of the $\Lambda$CDM model. This model is inspired by an exponential $f(R)$ gravity model studied recently in the literature [52, 53, 54], which depends upon two parameters, $b$ and $n$. Furthermore, instead of using the aforementioned strategies, we have chosen to implement the formalism developed in Ref. [55] within the context of $f(R)$ gravity. This formalism involves expressing the solution to the field equations, i.e., $H^{2}(N)$, as a perturbative series expansion in terms of the deviation parameter $b$ (making $n=1$), with the $\Lambda$CDM model solution as a reference. Additionally, we deduce approximated analytical expressions for the parameters $w_{\rm{DE}}$, $w_{\rm{eff}}$, and $\Omega_{\rm{DE}}$, as well as the deceleration parameter $q$ in terms of the redshift $z$. Finally, in order to verify the viability of this approximate analytical solution for $H(z)$ in late-times, we examine the behavior of these cosmological parameters, and perform a fit to observational data of the Hubble parameter ($H(z)_{{\rm obs}}$) and the apparent magnitude ($m_{b}(z)_{{\rm obs}}$, Pantheon SN sample) in terms of the redshift, to obtain constraints on the model parameter $b$, together with the present values of the expansion rate, $H_{0}$, and the matter density parameter, $\Omega_{m0}$.
@@ 2 The $f(Q)$ gravity: a brief review
$S=\int{d^{4}x\sqrt{-g}\left(\frac{f(Q)}{2\kappa^{2}}+\mathcal{L}_{\rm{M}}\right)},$ (2.1)
@@
$Q_{\gamma\mu\nu}\equiv\nabla_{\gamma}g_{\mu\nu}=\partial_{\gamma}g_{\mu\nu}-\Gamma^{\beta}_{\mkern 10.0mu\gamma\mu}g_{\beta\nu}-\Gamma^{\beta}_{\mkern 10.0mu\gamma\nu}g_{\beta\mu},$ (2.2)
@@
$\Gamma^{\lambda}\\ _{\mu\nu}=\left\{{}^{\lambda}\\ _{\mu\nu}\right\}+K^{\lambda}\\ _{\mu\nu}+L^{\lambda}\\ _{\mu\nu},$ (2.3)
@@
{}$\left\{{}^{\lambda}\\ _{\mu\nu}\right\}\equiv\frac{1}{2}g^{\lambda\beta}\left(\partial_{\mu}g_{\beta\nu}+\partial_{\nu}g_{\beta\mu}-\partial_{\beta}g_{\mu\nu}\right);$ (2.4)
@@
$K^{\lambda}\\ _{\mu\nu}\equiv\frac{1}{2}T^{\lambda}\\ _{\mu\nu}+T^{\,\,\,\,\,\lambda}_{(\mu\,\,\,\,\nu)},$ (2.5)
@@
$L^{\beta}_{\mkern 10.0mu\mu\nu}=\frac{1}{2}Q^{\beta}_{\mkern 10.0mu\mu\nu}-Q_{(\mu\nu)}^{\mkern 35.0mu\beta}.$ (2.6)
@@
$P^{\beta}_{\mkern 10.0mu\mu\nu}=-\frac{1}{2}L^{\beta}_{\mkern 10.0mu\mu\nu}+\frac{1}{4}(Q^{\beta}-\tilde{Q}^{\beta})g_{\mu\nu}-\frac{1}{4}\delta^{\beta}_{\,\,\,(\mu}Q_{\nu)},$ (2.8)
the other quantities present in Eq. (2.8) are given by
$Q_{\beta}=g^{\mu\nu}Q_{\beta\mu\nu},\quad\tilde{Q}_{\beta}=g^{\mu\nu}Q_{\mu\beta\nu}.$ (2.9)
Variation of the action (2.1) with respect to the metric gives the equation of motion [12, 13],
$\frac{2}{\sqrt{-g}}\nabla_{\beta}(f_{Q}\sqrt{-g}P^{\beta}_{\mkern 10.0mu\mu\nu})+\frac{1}{2}fg_{\mu\nu}+f_{Q}(P_{\mu\beta\lambda}Q_{\nu}^{\mkern 10.0mu\beta\lambda}-2Q_{\beta\lambda\mu}P^{\beta\lambda}\mkern 1.0mu\nu)=-T_{\mu\nu},$ (2.10)
@@
$T_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta\sqrt{-g}\mathcal{L}_{\rm{M}}}{\delta g^{\mu\nu}},$ (2.11)
@@
$ds^{2}=-dt^{2}+a(t)^{2}\delta_{ij}dx^{i}dx^{j},$ (2.12)
with $a(t)$ representing the scale factor, the time and spatial components of Eq. (2.10) are given, respectively, by [12, 13]
$6f_{Q}H^{2}-\frac{1}{2}f=\rho_{\rm{m}}+\rho_{\rm{r}},$ (2.13)
@@
$(12H^{2}f_{QQ}+f_{Q})\dot{H}=-\frac{1}{2}\left(\rho_{\rm{m}}+\frac{4}{3}\rho_{\rm{r}}\right),$ (2.14)
where $\rho_{\rm{m}}$ is the matter density and $\rho_{\rm{r}}$ denotes the density of radiation. $f_{QQ}=\frac{d^{2}f}{dQ^{2}}$, the over-dot denotes a derivative with respect to the cosmic time $t$ and $H\equiv\dot{a}/a$ is the Hubble parameter. Additionally, we have considered $\kappa^{2}=1$. Further, the non-metricity scalar associated with the metric (2.12) is given by
$Q=6H^{2}.$ (2.15)
@@
$\dot{\rho}_{\rm{m}}+3H\rho_{\rm{m}}=0,\quad\dot{\rho}_{\rm{r}}+4H\rho_{\rm{r}}=0.$ (2.16)
As usual in the literature, it is possible to rewrite the field equations (2.13) and (2.14) in the Einstein-Hilbert form:
$3H^{2}=\rho,$ (2.17)
$-2\dot{H}^{2}=\rho+p,$ (2.18)
where $\rho=\rho_{\rm{m}}+\rho_{\rm{r}}+\rho_{\rm{DE}}$ and $p=p_{\rm{m}}+p_{\rm{r}}+p_{\rm{DE}}$ correspond to the total effective energy pressure densities of the cosmological fluid, respectively. In this case, the dark energy component has a geometric origin, and after some manipulation in Eqs. (2.13) and (2.14), we obtain the effective dark energy and pressure corresponding to the $f(Q)$-theory given by
$\rho_{\rm{DE}}=\frac{1}{2}f+3H^{2}(1-2f_{Q}),$ (2.19)
@@
$w_{\rm{DE}}=\frac{p_{\rm{DE}}}{\rho_{\rm{DE}}}=-1+\frac{2\dot{H}(12H^{2}f_{QQ}+f_{Q}-1)}{\frac{1}{2}f+3H^{2}(1-2f_{Q})}.$ (2.22)
@@ 3 Cosmological dynamics in late-time
$f(Q)=Q+2\Lambda e^{-\left(\frac{b\Lambda}{Q}\right)^{n}},$ (3.1)
where $\Lambda$ is the cosmological constant, and $b$ and $n$ are real dimensionless parameters. This model is inspired in that studied in Refs. [52, 53, 54], but in the context of $f(R)$ gravity. It is evident that for $b=0$ the model given by Eq. (3.1) is equivalent to GR plus the cosmological constant. In particular, from the structure of this model, it can be seen as a perturbative deviation from the $\Lambda$CDM Lagrangian. In this sense, this model can be arbitrarily close to $\Lambda$CDM, depending on the parameter $b$, i.e., the exponential form presented in the Eq. (3.1) elegantly and succinctly encapsulates the perturbative nature of the model across all orders in the parameter $b$, which is evident from a Taylor series expansion about this parameter,
$f(Q)=(Q+2\Lambda)-\frac{2\Lambda^{2}b}{Q}+\frac{\Lambda^{3}b^{2}}{Q^{2}}-\frac{\Lambda^{4}b^{3}}{3Q^{3}}+\frac{\Lambda^{5}b^{4}}{12Q^{4}}-\cdots,$ (3.2)
where for simplicity we have considered $n=1$. It should be highlighted that in the literature also other exponential $f(Q)$ gravity models have been intensively studied (see e.g., Refs. [22, 28, 29, 30, 36, 37, 38, 41]).
Following the procedure carried out in Ref. [55], we rewrite Eq. (2.13) in terms of $N=\ln{a}$
$2f_{Q}H^{2}(N)-\frac{1}{6}f-(\Omega_{m0}e^{-3N}+\Omega_{r0}e^{-4N})H_{0}^{2}=0,$ (3.3)
where we have introduced the current energy density parameters, $\Omega_{m0}=\frac{\rho_{m0}}{3M_{\rm{p}}^{2}H_{0}^{2}}$ and $\Omega_{r0}=\frac{\rho_{r0}}{3M_{\rm{p}}^{2}H_{0}^{2}}$; here $M_{\rm{p}}=1$, and $H_{0}$ corresponds to the current Hubble parameter. Implicitly, we consider by simplicity $n=1$ and $\Lambda=3(1-\Omega_{m0}-\Omega_{r0})H_{0}^{2}$. Replacing Eq. (3.1) in Eq. (3.3) and taking into account Eq. (2.15), we obtain an algebraic equation for $H^{2}(N)$, given by
$H^{2}(N)-(\Omega_{m0}e^{-3N}+\Omega_{r0}e^{-4N})H_{0}^{2}$ (3.4)
@@
Since the above is a transcendental equation, it cannot be solved for $H^{2}(N)$ in explicit form. Another possibility involves constructing a first-order differential equation for $H^{2}(N)$ using Eqs. (2.13) and (2.14). However, in this scenario the equation is nonlinear and lacks an analytical solution for $H^{2}(N)$.
Usually, in the literature for circumventing this issue, a common practice is to propose a parametrization of the Hubble parameter in terms of the redshift [17, 31, 43], among other strategies. In this work instead of use that strategy, we use the formalism developed in Ref. [55] in the context of $f(R)$ gravity, which consists in expressing the solution to Eq. (3.3), i.e., $H^{2}(N)$, as a perturbative series expansion in terms of the deviation parameter $b$, with respect to the $\Lambda$CDM model solution as a reference, as follows
$H^{2}(N)=H^{2}_{\Lambda}(N)+\sum_{i=1}^{M}b^{i}\delta H_{i}^{2}(N),$ (3.5)
@@
$\frac{H^{2}_{\Lambda}(N)}{H_{0}^{2}}=\Omega_{m0}\,e^{-3N}+\Omega_{r0}\,e^{-4N}+(1-\Omega_{m0}-\Omega_{r0})=E_{\Lambda}^{2}(N),$ (3.6)
further, as has been demonstrate in Ref. [55], we can consider only two terms in the above series expansion; in doing so, Eq. (3.5) reduces to
$H^{2}(N)\approx H^{2}_{\Lambda}(N)+b\,\delta H_{1}^{2}(N)+b^{2}\delta H_{2}^{2}(N).$ (3.7)
Replacing Eq. (3.7) in Eq. (3.4) and considering up to second order in $b$, it is straightforward to calculate $\delta H_{1}^{2}(N)$ and $\delta H_{2}^{2}(N)$, which are given by
$\frac{\delta H_{1}^{2}(N)}{H_{0}^{2}}=-\frac{3H_{0}^{2}(\Omega_{m0}+\Omega_{r0}-1)^{2}}{2H_{\Lambda}^{2}(N)},$ (3.8)
$\frac{\delta H_{2}^{2}(N)}{H_{0}^{2}}=-\frac{H_{0}^{4}(\Omega_{m0}+\Omega_{r0}-1)^{3}[18H_{0}^{2}(\Omega_{m0}+\Omega_{r0}-1)+5H_{\Lambda}^{2}(N)]}{8(H_{\Lambda}^{2}(N))^{3}}.$ (3.9)
Now, replacing Eqs. (3.8) and (3.9) in Eq. (3.7), and using Eq. (3.6), we obtain an approximate solution for the Hubble parameter $H(z)$:
$E^{2}(z)=$ $\frac{H^{2}(z)}{H_{0}^{2}}\approx 1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}-\frac{3(\Omega_{m0}-1)^{2}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]}b$ (3.10)
@@
$w_{\rm{DE}}=-1+\frac{1}{3}(1+z)\frac{(\rho_{\rm{DE}}(z))^{\prime}}{\rho_{\rm{DE}}(z)},$ (3.11)
$w_{\rm{eff}}=-1+\frac{1}{3}(1+z)\frac{(E^{2}(z))^{\prime}}{E^{2}(z)},$ (3.12)
$\Omega_{\rm{DE}}=1-\frac{\Omega_{m0}(1+z)^{3}}{E^{2}(z)};$ (3.13)
@@
$q=-1+\frac{1}{2}(1+z)\frac{(E^{2}(z))^{\prime}}{E^{2}(z)},$ (3.14)
where the prime denotes differentiation with respect to $z$. Using Eqs. (2.19) and (3.10) and considering up to second order expansion in $b$, we obtain approximated analytical expressions for the above parameters in terms of the redshift $z$, as follows:
$w_{\rm{DE}}\approx$ $-1-\frac{3[(1+z)^{3}(\Omega_{m0}-1)\Omega_{m0}]}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{2}}b$ (3.15)
@@
$w_{\rm{eff}}\approx$ $-1+\frac{(1+z)^{3}\Omega_{m0}}{1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}}+\frac{3(1+z)^{3}(\Omega_{m0}-1)^{2}\Omega_{m0}}{[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{3}}b$ (3.16)
@@
$\Omega_{\rm{DE}}\approx$ $1-\frac{(1+z)^{3}\Omega_{m0}}{1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}}-\frac{3(1+z)^{3}(1-\Omega_{m0})^{2}\Omega_{m0}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{3}}b$ (3.17)
@@
$q\approx$ $-1+\frac{3(1+z)^{3}\Omega_{m0}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]}+\frac{9(1+z)^{3}(\Omega_{m0}-1)^{2}\Omega_{m0}}{2[1-\Omega_{m0}+(1+z)^{3}\Omega_{m0}]^{3}}b$ (3.18)
@@
With Eqs. (3.15)-(3.18), we can plot the evolution of each parameter in terms of the redshift $z$. Additionally, in order to compare the results with the $\Lambda$CDM model, we have also incorporated in the corresponding plots the behavior associated with each quantity defined by Eqs. (3.11)-(3.14), but using Eq. (3.6) instead of (3.10).